OpenStudy (anonymous):

g(x)=-x+4 f(x)=-2(abs(x-2))+4 Find g(f(x)) g(x)=-x+4 f(x)=-2(abs(x-2))+4 Find g(f(x)) @Mathematics

hero (hero):

g(f(x)) = -2|x-2|+4+4 = -2|x-2|+8

OpenStudy (anonymous):

Oh cool that is the answer I got. Thanks hero

hero (hero):

Well, actually

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

Oh ya i see that

hero (hero):

There's a negative in front of x

hero (hero):

-(-2|x-2|+4)+4 Write it like that instead

OpenStudy (anonymous):

But lets say I have to find the derivative it would be easier to open the brackets

hero (hero):

Well, the absolute value kind of complicates things

OpenStudy (anonymous):

Lets say the derivative of v'(1)

hero (hero):

Well, if that's the case, then it is probably zero

OpenStudy (anonymous):

alright I will use the site you gave me yesterday but I don't know how I will show absolute value

hero (hero):

once you replace x with a number, and evaluate it, then the derivative of any number is just zero

OpenStudy (anonymous):

Oh so maybe I should defferentiate it first and then replace the x value with 1

OpenStudy (anonymous):

hold the phone. it is true that the derivative of a constant is zero. but that means the derivative of a constant function is 0

OpenStudy (anonymous):

so for example the derivative of 4 is 0, but the derivative of \[f(x)=x^2\] at x = 4 is \[f'(4)=2\times 4=8\] not 0!

OpenStudy (anonymous):

ya true

OpenStudy (anonymous):

so if you want \[v'(1) \] the first thing you need is \[v'(x)\]

OpenStudy (anonymous):

and for that the first thing you need is \[v(x)\] without absolute values signs

OpenStudy (anonymous):

so you have \[v(x)=4|x-2|+8\] i believe. and what this is depends on whether x is greater than or less than 2

hero (hero):

Where'd you get the 4 from?

OpenStudy (anonymous):

Oh ya? that is cool!!!!

OpenStudy (anonymous):

if x is greater than 2 you get \[v(x)=2(x-2)+8=2x\] whereas if x is less than 2 you get \[v(x)=2(2-x)+8=12-2x\]

OpenStudy (anonymous):

damn two typos

OpenStudy (anonymous):

@hero the 4 is a straight up mistake. it is 2, not 4

OpenStudy (anonymous):

and also \[2(x-2)+8=2x+4\]

hero (hero):

Yeah, but like you said, it's not the only one

OpenStudy (anonymous):

in any case since 1 is less than 2, use the second formula, and see that the derivative is -2

OpenStudy (anonymous):

oh great! Thanks for your help

hero (hero):

There's still another mistake lurking

hero (hero):

If I'm not mistaken, distribution comes before addition

OpenStudy (anonymous):

yes it does!!!

hero (hero):

v(x) = -(-2|x-2|+4)+4

hero (hero):

I'm wondering if that simplifies to v(x) = 2|x-2|

OpenStudy (anonymous):

I don't know

hero (hero):

I guess satelliate gave up

OpenStudy (anonymous):

I guess so ;D

hero (hero):

If we continue with his logic, it splits into two: v(x) = 2(x-2) v(x) = 2(2-x)

hero (hero):

v(x) = 2x-4 v(x) = 4 - 2x v'(x) = 2 v'(x) = -2

hero (hero):

I don't know which one we use though

OpenStudy (anonymous):

I will play aound with it and see what I get

OpenStudy (anonymous):

It is -2 I just calculated it on my graphing calc

hero (hero):

Good

OpenStudy (anonymous):

But i am confused as to how you got the equation of the derivative

OpenStudy (anonymous):

Did you use the chain rule

hero (hero):

Yes, why?

hero (hero):

You don't know chain rule yet?

OpenStudy (anonymous):

Ya i do

OpenStudy (anonymous):

Hey look the derivative we got was totally wrong! That is interesting

hero (hero):

Oh well. At least you found the right one

hero (hero):

Actually, it's still 2

hero (hero):

Take a closer look

hero (hero):

sqrt and squared cancel in denominator (x-2) cancels top and bottom

hero (hero):

:D

OpenStudy (anonymous):

Oh I will check that out

OpenStudy (anonymous):

i am sure you are long gone, but just for the record that is not 2. it is 2 if x is greater than 2, -2 if x is less than 2. just thought i would mention it.

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

Oh I will check that out

OpenStudy (anonymous):

hero wrote out the correct answer above he wrote v(x) = 2x-4 v(x) = 4 - 2x v'(x) = 2 v'(x) = -2

OpenStudy (anonymous):

the only thing missing is for which values of x you use the different formulas. it is v(x) = 2x-4 if x > 2 v(x) = 4 - 2x if x < 2 v'(x) = 2 if x > 2 v'(x) = -2 if x < 2

OpenStudy (anonymous):

and therefore \[v'(1)=-2\] whereas \[v'(5)=2\]

OpenStudy (anonymous):

I graphed the derivative and it only shows a straight line starting at x=2 until positive infinity at y=2

OpenStudy (anonymous):

Ya you are correct

OpenStudy (anonymous):

the graph of the derivative is two horizontal lines

OpenStudy (anonymous):

oh great thanks